Circular No. 7375

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000O IN MCG +3-31-61 T. Puckett, Mountain Town, GA, reports the discovery of an appar-

ent supernova (mag 17.1) on an unfiltered CCD frame (limiting mag 19.2) taken with the Puckett Observatory 0.35-m automated supernova patrol telescope on Mar. 5.26 UT. The candidate is located at α = 12^{h}11^{m}15^{s}.58, δ = +17^{o}53′32′′.2 (equinox 2000.0), which is 7′′.7 east and 14′′.6 north of the center of MCG +3-31-61. The new object was also present on an unfiltered CCD frame taken on Mar. 6.10, but it was not definitely present (though there was possibly a hint) on a frame taken on Feb. 4.32, neither does it appear on Palomar Sky Survey images taken on 1990 Mar. 23 (limiting mag ∼ 21.0) and 1950 Apr. 20 (limiting mag ∼ 20.0).

SUPERNOVA 2000H IN IC 454 S. Benetti, E. Cappellaro, M. Turatto, and A. Pastorello, Padua, re-

port: “We have obtained two good S/N spectra of SN 2000H (cf. IAUC 7366, 7367) on Mar. 2.0 and 3.0 UT with the European Southern Observatory 3.6-m telescope (+ EFOSC2). The spectra (range 340–1020 nm, resolution 1.5 nm) are now dominated by P-Cyg lines of He I (447.1, 501.5, 587.6, 667.8, 706.5 nm), from which a mean expansion velocity of 8000 ± 1000 km/s is derived. In the red part of the spectrum, the Ca II infrared triplet is present and intense, from which an expansion velocity of ∼ 7600 km/s is derived. At 637.6 nm, there is a deep absorption that we identify with Hα and from which an expansion velocity of ∼ 12 500 km/s is derived. In the blue part of the spectrum, there is a tiny absorption centered at 472.0 nm, from which (if identified with Hβ) an expansion velocity of 12 400 km/s is obtained, so strengthening the hypothesis of the hydrogen presence in the SN 2000H. Given the simultaneous presence of strong He and H, based on these spectra, the supernova should be probably classified as of type IIB (a core collapse with very little H left in the ejecta), possibly similar to SN 1993J. We note also a narrow absorption at 596.6 nm that we attribute to a Na D I interstellar line originating in the parent galaxy. The relatively large equivalent width (0.12 nm) suggests that the supernova suffered a reddening of E(B − V) ∼ 0.16 (Benetti et al., in preparation).”

SUPERNOVA 1999gi IN NGC 3184 Unfiltered CCD magnitudes by S. Yoshida and K. Kadota, Ageo, Sai-

tama, Japan: 1999 Dec. 19.671 UT, 14.5; 22.657, 14.5; 29.650, 14.5; 2000 Jan. 28.650, 14.5; Feb. 12.483, 14.6; 27.690, 14.7.

2000 March 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green